OpenStudy (anonymous):

An aircraft has a mass of 3.60 x 105 kg. At a certain instant during its landing, its speed is 50.6 m/s. If the braking force is 3.20 x 105 N, what is the speed of the airplane 19.4 s later?

OpenStudy (anonymous):

force =mass*acceleration -3.2*105 N = (3.6*105)*a (the force is negative because it's in the opposite direction of the motion of the aircraft) solve for a a=-.88889 (acceleration is negative because the aircraft is slowing down) velocity = initial velocity plus acceleration times time v= vi+at we know vi=50.6 m/s and a=.88889 and t=19.4 so now just plug in all the numbers to get v=33.3556 m/s